# FIPER: Generalizable Factorized Fields for Joint Image Compression and Super-Resolution

### v1

@@ Abstract
This paper introduces a framework that integrates image compression and super-resolution using a unified representation called Generalizable Factor Fields. Traditional approaches to image compression often lead to significant information loss, particularly at low bitrates, which blocks image usability for both human viewing and computer vision applications. Super-resolution techniques strive to reconstruct high-quality images from these degraded compressions but face challenges when starting with heavily compressed sources. Our approach leverages a coefficient-basis structure within the Factor Fields, optimizing the joint task of compressing multiple images and enhancing their resolution in a single cohesive process. By encoding and transmitting images using shared basis functions, our method significantly reduces information loss and boosts efficiency in handling large image datasets. We demonstrate superior performance in both super-resolution and compression tasks compared to existing methods, offering improved image reconstruction quality and more efficient low-bitrate operation.
@@ 1 Introduction
Single Image Super-Resolution (SISR) aims to reconstruct high-quality images from low-resolution counterparts. Specifically, the key lies in accurately restoring fine details and reconstructing the correct arrangement of visual features. Thus, geometric correspondences or repetitive patterns, such as stripes, grids, or textures, are commonly used for evaluation due to their rich details that are crucial to image fidelity. Early CNN-based approaches flickr2k, SRCNNlaid the foundation for SISR, which was later enhanced with GAN-based methods srgan for improved perceptual realism. Follow-up Transformer-based networks IPT address non-local dependencies, with subsequent Swin Transformer-based approaches sparking tremendous advancements swin2sr, DAT, RGT, which then inspire more and more complex, delicate designs of heavy network architectures in SISR. However, these prior works have primarily focused on network architecture design rather than addressing the capability of representation. The visual patterns and the inherent nature of image content structure that play an influential role in SISR have not been explicitly considered in the representation learning process. This raises a critical question: beyond a simple network output, can we derive a formulation that more effectively captures these patterns and aligns with the goals of SISR?
@@
On the other hand, image compression serves as a fundamental task in low-level vision applications, where the traditional compression standards VVC-21.2, jpeg, jpeg2000 lay the groundwork. The emerging learned image compression models imc_cnn1, ballé2018variational, guo-hua2023evc, imc_lic_tcm, minnen2018joint, cheng2020learned, compression algorithms of which mostly follow the pixel-space transform coding  imc_cnn1, Goyal2001paradigm, then introduce neural networks to further optimize compression efficiency by learning more compact latent representations and improving reconstruction quality. Specifically, they convert pixels into compact representations through a transform module, which eliminates the redundancy and reduces the bit cost in the subsequent entropy coding process. However, the core challenge of image compression is to accurately reconstruct the information lost during compression and quantization. In other words, the models are to recover the impaired components in images, which can essentially be viewed as reconstructing a high-quality image from its ‘low-resolution’ version, much like Super-Resolution.
Based on the aforementioned analysis, it becomes clear that although Super-Resolution and Image Compression appear to be two distinct tasks, they share mutual similarities in two key aspects: (1) Both tasks require models to restore fine details from low-quality image content, as well as implicitly capture and reconstruct repetitive structural elements. (2) Both aim to conserve image quality, either by enhancing resolution or efficiently compressing data without significant loss of perceptual fidelity.
@@
Hence, inspired by recent advances in decomposition fields and matrices factorization in 3D scene modeling tensorf, mueller2022instantngp, eg3d, kplanes, cao2023hexplane, factor-field, gao2023strivec, we propose a unified representation, Factorized Fields, with generalizable Coefficient Backbone and Basis Transformer. This approach explicitly captures multi-scale visual features and repetitive structural components in images through a basis-coefficient decomposition. The resulting representation strikes a balance between being compact and information-rich, enabling the resolution of structural ambiguities and the precise modeling of image details through a multi-frequency formulation. In the meantime, such a formulation imposes a factorization constraint during model training, which not only enhances the quality of single-image super-resolution (SISR) but also reduces distortion and improves compression efficiency by explicitly modeling structural elements. On top of these, to leverage the robust information-recovering capability of SR models, we integrate such priors with Image Compression models, whose knowledge of detail compensation further refines the lost key elements.
@@
  • We propose Factorized Fields, a unified representation that explicitly models multi-scale visual features and structural components for both super-resolution and image compression.
@@ 2 Related Works
Image super-resolution is critical in computer vision, focusing on recovering high-resolution (HR) images from low-resolution (LR) inputs. Following the foundational studies, CNN-based strategies SRCNN, flickr2k, zhou2020cross, shufflemixer, VDSR were initially introduced with modeling techniques such as residual learning ledig2017photo, liu2020residual, tong2017image, zhang2018residual, RCAN, drct, EDSR , or recursive learning RGT, tai2017recursive. Besides, subsequent research also sheds light on GAN-base methods srgan, ledig2017photo, esrgan, real_esrgan to enhance realisticity and detail quality. However, the inductive bias of CNN-based networks by restricting spatial locality hinders the capture of long-range dependency from images, which is alleviated by Transformer-based SISR networks IPT, EDT. Afterward, SwinIR swinir is proposed to combine spatial locality and non-local information by Swin Transformer liu2021swin with window attention and achieve breakthrough improvement in SISR. Following SwinIR’s success, several works have built upon its framework swin2sr, zhu2023attention, ATD to reach better image quality as well as solve information bottleneck drct. Hybrid approaches CRAFT (CRAFT) merge the benefits of convolutional and transformer structures to further elevate SR performance. For better feature aggregation, DAT (DAT) and HAT (chen2023hat) integrate spatial and channel information using attention mechanisms to enhance their representation capabilities. Moreover, RGT (RGT) introduces a unique recursive-generalization self-attention mechanism that efficiently captures comprehensive spatial details with a linear increase in computational complexity.
@@
Deep learning has significantly advanced image compression, offering superior compression ratios and image quality compared to traditional methods like JPEG jpeg and JPEG-2000 jpeg2000. Early CNN-based approaches imc_cnn1, imc_cnn2, imc_cnn3 have been surpassed by transformer-based models imc_trans1, imc_trans2, imc_trans3, duan2023cross-component, which leverage spatio-channel attention for better performance. GAN-based methods imc_gan1, imc_gan2, imc_gan3, imc_gan4 have further contributed to real-time adaptive compression. Recently, ELIC imc_elic introduced efficient compression with unevenly grouped space-channel contextual adaptive coding, while LIC TCM imc_lic_tcm integrated transformers and CNNs to capture both global and local image features. The eContextformer koyuncu2024efficient introduced patch-wise, checkered, and channel-wise grouping techniques for parallel context modeling with a shifted window spatial-channel attention mechanism. GroupedMixer li2024groupedmixer proposed a transformer-based entropy model with token mixers for inner and cross-group context modeling. Meanwhile, Wavelet Conditional Diffusion song2024wavelet-compression introduced a wavelet-based model with uncertainty-aware loss, balancing high perceptual quality with low distortion.
@@ 3 Methods
In this section, we first briefly introduce the background of the factor fields factor-field, their properties, and also the preliminary of Learned Image Compression in Sec. 3.1. Inspired by factor fields, we next explain the motivation and how we derive our Factorized Fields for enhanced image reconstruction quality in Sec. 3.2. We then describe how to adapt the formulation to Super-Resolution in Sec. 3.3. Finally, we show how to incorporate such representation with image(s) compression and how we integrate Super-Resolution and Image Compression in Sec. 3.4.
Figure 1: The correlation between coordinate transformation and downsampling. (a) The sawtooth transformation example with $k=2$. (b) The PixelUnShuffle downsample. (c) To explicitly model the information for sampling with a sawtooth, we rearrange the feature map in a dilation-like manner in the downsample layer of the Basis Swin Transformer. This way, the feature sampled would capture the information in the original layout correctly.
3.1 Preliminaries
Factor Fields.
The concept of decomposition fields or factorized matrices in the reconstruction of 2D images or 3D scenes has shown superior rendering quality and improved efficiency recently tensorf, gao2023strivec, mueller2022instantngp, kplanes, cao2023hexplane, eg3d, while factor-field first proposed a unified representation called factor fields. For a 1D signal $s(x)$, the factor fields formulation is:
$\hat{s}(x)=c(x)^{T}b(\gamma(x)),$ (1)
where $c(x)=(c_{1}(x),...,c_{N}(x))^{T}$ are spatial-varying coefficient fields, $b(x)=(b_{1}(x),...,b_{N}(x))^{T}$ are basis functions, and $\gamma(x)$ is a coordinate transformation. The coefficient and basis are queried by sampling with the coordinates $x$. This generalizes to Q-dimensional signals:
$\hat{s}(x)=\mathcal{P}(c(x)\odot b(\gamma(x))),$ (2)
where $\mathcal{P}$ is a projection function and $\odot$ denotes the element-wise product. In practice, for multi-dimensional coefficients:
@@ 3.1 Preliminaries
This formulation allows the same basis to be applied at multiple spatial locations while varying the coefficients, especially when $\gamma$ is a periodic function.
Learned Image Compression.
Following (channel-autoregressive, ballé2018variational), a learned image compression model with a channel-wise entropy model can be formulated as:
$z$ $=h_{a}(y;\phi_{h}),\ y=g_{a}(x;\phi),$ (4)
$\{X_{\text{mean}},X_{\text{scale}}\}$ $=h_{s}(\hat{z};\theta_{h}),\ \hat{z}=Q(z)$
$\hat{y}$ =$=\{Q(y_{0}-\mu_{0})+\mu_{0},...,Q(y_{t}-\mu_{t})+\mu_{t}\},\ 0<=t<l,\ \mu_{t}=e_{i}(\hat{y}_{<i},X_{\text{mean}}))$
$\hat{x}$ $=g_{s}(\overline{y};\theta),\ \overline{y}=\textbf{Refine}_{\theta_{r}}({\mu_{0},...,\mu_{t}},\hat{y}).$
The encoder $g_{a}$ transforms the raw image $x$ into a latent representation $y$. A hyper-prior encoder $h_{a}$ further processes $y$ to output $z$, capturing spatial dependencies. $z$ is quantized to $\hat{z}$, which is decoded by $h_{s}$ to produce features $X_{\text{mean}}$ and $X_{\text{scale}}$, used to estimate the mean $\mu$ and variance $\sigma$ of $y$. The latent $y$ is divided into $l$ slices, and each quantized around computed means $\mu_{t}$. These means are derived from earlier quantized slices and $X_{\text{mean}}$ by a slice network $e_{i}$. The quantized slices form $\hat{y}$. For decompression, $\hat{y}$ is refined using $\textbf{Refine}_{\theta_{r}}$ based on $\mu_{t}$ and $\hat{y}$ to produce $\overline{y}$, approximating the original $y$. Finally, $g_{s}$ reconstructs the decompressed image $\hat{x}$ from $\overline{y}$. The model is trained using a Lagrangian multiplier-based rate-distortion optimization:
$L=R(\hat{y})+R(\hat{z})+\lambda\cdot D(x,\hat{x}),$ (5)
where $R(\hat{y})$ and $R(\hat{z})$ denote bit rates, $D(x,\hat{x})$ is the distortion term (calculated by MSE), and $\lambda$ balances compression efficiency and image fidelity. In our experiments, we follow imc_lic_tcm, modifying only $g_{s}$ to demonstrate our representation’s effectiveness.
@@ 3.2 Formulation of Factorized Fields
As discussed in Sec. 1, the key to superior rendering quality in image regression tasks of both Image Compression and Super-Resolution lies in the capability of the representation to capture accurate structural distribution and fine visual details. Meanwhile, transformations such as Fourier Transform or Wavelet Transform have long been used to model multi-frequency information fuoli2021fourier, waveletsr to express different implicit functions in images; however, such methods often suffer from under-expression due to the pre-defined and limited frequency bands or restriction from its formulation. Thus, we seek a representation that explicitly incorporates and fits multi-scale and multi-frequency components and yet is highly flexible and learnable according to individual features in an image.
Inspired from recent success on such decomposition fields kplanes, factor-field, tensorf, mueller2022instantngp, cao2023hexplane, eg3d, we primarily built our Factorized Fields framework on factor fields factor-field from Eq. 3:
@@
Such formulation has several key properties: First of all, by decomposing the images into basis frequencies, we can learn the implicit functions of an image and capture the mutual dependencies between pixels and across spatial composition; meanwhile, since the basis and coefficient are specific to every single image and both learnable in all spatial dimensions, the restriction on a limited number of basis (we use $N=6$ in all of our experiments) can be alleviated. Finally, as in factor-field, of all the tested coordinate transformation $\gamma$ in Eq. 3 sawtooth $\gamma(x)=x\bmod{k},k\in\mathbb{R}$ performs mostly the best in image regression tasks. We can easily observe that such transformation implicitly captures patch-like frequency information as shown in Fig. 1a, and thus, we propose that by leveraging the inter-patch information from the sawtooth coordinate transformation, the visual correspondence between spatial locations can be effectively represented.
However, in practice, we sample the basis via bilinear or bicubic sampling due to memory constraints, i.e., the feature size is less than image height and width, and this poses a severe problem: These nearby sampled pixel features are actually linear or based on cubic interpolation with respect to the basis, and such inductive bias hinders the representation of non-linearity in images. To resolve this, we propose a modified version, our Factorized Fields:
$\hat{I}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{i}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},\psi\in\{\text{sin},\text{cos}\},\alpha_{j}\in R.$ (7)
Here, inspired by the Fourier Series, with a scalar $\alpha$ which can be viewed as increasing to sampling frequency and a transformation function $\psi$ to add to implicit non-linearity functions and modulation of product, our complete Factorized Fields method stands out by the fact that such formulation significantly enhances the ability to represent complex non-linear structures in images and effectively composes high-frequency components between pixels.

### v2

@@ Abstract
In this work, we propose a unified representation for Super-Resolution (SR) and Image Compression, termed Factorized Fields, motivated by the shared principles between these two tasks. Both SISR and Image Compression require recovering and preserving fine image details—whether by enhancing resolution or reconstructing compressed data. Unlike previous methods that mainly focus on network architecture, our proposed approach utilizes a basis-coefficient decomposition to explicitly capture multi-scale visual features and structural components in images, addressing the core challenges of both tasks. We first derive our SR model, which includes a Coefficient Backbone and Basis Swin Transformer for generalizable Factorized Fields. Then, to further unify these two tasks, we leverage the strong information-recovery capabilities of the trained SR modules as priors in the compression pipeline, improving both compression efficiency and detail reconstruction. Additionally, we introduce a merged-basis compression branch that consolidates shared structures, further optimizing the compression process. Extensive experiments show that our unified representation delivers state-of-the-art performance, achieving an average relative improvement of 204.4% in PSNR over the baseline in Super-Resolution (SR) and 9.35% BD-rate reduction in Image Compression compared to the previous SOTA.
@@ 1 Introduction
Figure 1: The Proposed Factorized Fields. With explicit formulation of repetitive patterns and high frequency components, images can be represented with fine-grained details, especially useful for both Super-Resolution and Image Compression. Note that in this figure $\alpha=\{1,16\}$ and $\psi=sin$.
@@
Single Image Super-Resolution (SISR) aims to reconstruct high-quality images from low-resolution counterparts. Specifically, the key lies in accurately restoring fine details and reconstructing the correct arrangement of visual features. Thus, geometric correspondences or repetitive patterns, such as stripes, grids, or textures, are commonly used for evaluation due to their rich details that are crucial to image fidelity. Early CNN-based approaches [46, 21]laid the foundation for SISR, which was later enhanced with GAN-based methods [39] for improved perceptual realism. Follow-up Transformer-based networks [12] address non-local dependencies, with subsequent Swin Transformer-based approaches sparking tremendous advancements [19, 15, 16], which then inspire more and more complex, delicate designs of heavy network architectures in SISR. However, these prior works have primarily focused on network architecture design rather than addressing the capability of representation. The visual patterns and the inherent nature of image content structure that play an influential role in SISR have not been explicitly considered in the representation learning process. This raises a critical question: beyond a simple network output, can we derive a formulation that more effectively captures these patterns and aligns with the goals of SISR?
@@
On the other hand, image compression serves as a fundamental task in low-level vision applications, where the traditional compression standards [33, 69, 66] lay the groundwork. The emerging learned image compression models [7, 2, 29, 48, 58, 17], compression algorithms of which mostly follow the pixel-space transform coding  [7, 27]paradigm, then introduce neural networks to further optimize compression efficiency by learning more compact latent representations and improving reconstruction quality. Specifically, they convert pixels into compact representations through a transform module, which eliminates the redundancy and reduces the bit cost in the subsequent entropy coding process. However, the core challenge of image compression is to accurately reconstruct the information lost during compression and quantization. In other words, the models are to recover the impaired components in images, which can essentially be viewed as reconstructing a high-quality image from its ‘low-resolution’ version, much like Super-Resolution.
@@
Based on the aforementioned analysis, it becomes clear that although Super-Resolution and Image Compression appear to be two distinct tasks, they share mutual similarities in two key aspects: (1) Both tasks require models to restore fine details from low-quality image content, as well as implicitly capture and reconstruct repetitive structural elements. (2) Both aim to conserve image quality, either by enhancing resolution or efficiently compressing data without significant loss of perceptual fidelity. Hence, inspired by recent advances in decomposition fields and matrices factorization in 3D scene modeling [9, 59, 8, 24, 6, 10, 26], we propose a unified representation, Factorized Fields, with generalizable Coefficient Backbone and Basis Transformer. This approach explicitly captures multi-scale visual features and repetitive structural components in images through a basis-coefficient decomposition. The resulting representation strikes a balance between being compact and information-rich, enabling the resolution of structural ambiguities and the precise modeling of image details through a multi-frequency formulation. In the meantime, such a formulation imposes a factorization constraint during model training, which not only enhances the quality of single-image super-resolution (SISR) but also reduces distortion and improves compression efficiency by explicitly modeling structural elements. On top of these, to leverage the robust information-recovering capability of SR models, we integrate such priors with Image Compression models, whose knowledge of detail compensation further refines the lost key elements.
@@
  • We propose Generalized Factorized Fields, a unified representation that explicitly models multi-scale visual features and structural components for both super-resolution and image compression.
@@ 2 Related Works
Image super-resolution is critical in computer vision, focusing on recovering high-resolution (HR) images from low-resolution (LR) inputs. Following the foundational studies, CNN-based strategies [21, 46, 79, 64, 34] were initially introduced with modeling techniques such as residual learning [40, 47, 68, 78, 77, 18, 45] , or recursive learning [16, 65]. Besides, subsequent research also sheds light on GAN-base methods [39, 40, 70, 71] to enhance realisticity and detail quality. However, the inductive bias of CNN-based networks by restricting spatial locality hinders the capture of long-range dependency from images, which is alleviated by Transformer-based SISR networks [12, 43]. Afterward, SwinIR [44] is proposed to combine spatial locality and non-local information by Swin Transformer [50] with window attention and achieve breakthrough improvement in SISR. Following SwinIR’s success, several works have built upon its framework [19, 80, 75] to reach better image quality as well as solve information bottleneck [18]. Hybrid approaches CRAFT [41] merge the benefits of convolutional and transformer structures to further elevate SR performance. For better feature aggregation, DAT [15] and HAT [13] integrate spatial and channel information using attention mechanisms to enhance their representation capabilities. Moreover, RGT [16] introduces a unique recursive-generalization self-attention mechanism that efficiently captures comprehensive spatial details with a linear increase in computational complexity.
@@
Deep learning has significantly advanced image compression, offering superior compression ratios and image quality compared to traditional methods like JPEG [69] and JPEG-2000 [66]. Early CNN-based approaches [7, 3, 28] have been surpassed by transformer-based models [37, 81, 72, 22], which leverage spatio-channel attention for better performance. GAN-based methods [53, 54, 23, 62] have further contributed to real-time adaptive compression. Recently, ELIC [30] introduced efficient compression with unevenly grouped space-channel contextual adaptive coding, while LIC TCM [48] integrated transformers and CNNs to capture both global and local image features. The eContextformer [38] introduced patch-wise, checkered, and channel-wise grouping techniques for parallel context modeling with a shifted window spatial-channel attention mechanism. GroupedMixer [42] proposed a transformer-based entropy model with token mixers for inner and cross-group context modeling. Meanwhile, Wavelet Conditional Diffusion [63] introduced a wavelet-based model with uncertainty-aware loss, balancing high perceptual quality with low distortion.
@@ 3 Methods
In this section, we present the details of our proposed methods. Inspired by factor fields, which can be formulated by
@@
and whose details can be found in supplementary materials , we first explain the motivation and how we derive our Factorized Fields for enhanced image reconstruction quality in  Sec. 3.1. We then describe how to adapt the formulation to Super-Resolution in  Sec. 3.2. Finally, we show how to incorporate such representation with image(s) compression and how we integrate Super-Resolution and Image Compression in  Sec. 3.3. Note that the background of learned image compression is described in supplementary materials.
Figure 2: The correlation between coordinate transformation and downsampling. (a) The sawtooth transformation example with $k=2$. (b) The PixelUnShuffle downsample. (c) To explicitly model the information for sampling with a sawtooth, we rearrange the feature map in a dilation-like manner in the downsample layer of the Basis Swin Transformer. This way, the feature sampled would capture the information in the original layout correctly; in other words, $I=\text{Sawtooth-Pixel-Sampling}(\text{Sawtooth-Aware-Downsample}(I))$
@@ 3.1 Formulation of Factorized Fields
As discussed in  Sec. 1, the key to superior rendering quality in image regression tasks of both Image Compression and Super-Resolution lies in the capability of the representation to capture accurate structural distribution and fine visual details. Meanwhile, transformations such as Fourier Transform or Wavelet Transform have long been used to model multi-frequency information [25, 36] to express different implicit functions in images; however, such methods often suffer from under-expression due to the pre-defined and limited frequency bands or restriction from its formulation. Thus, we seek a representation that explicitly incorporates and fits multi-scale and multi-frequency components and yet is highly flexible and learnable according to individual features in an image.
Inspired from recent success on such decomposition fields [24, 10, 9, 59, 6, 8], we primarily built our Factorized Fields framework on factor fields [10] from  Eq. 1:
@@
Such formulation has several key properties: First of all, by decomposing the images into basis frequencies, we can learn the implicit functions of an image and capture the mutual dependencies between pixels and across spatial composition; meanwhile, since the basis and coefficient are specific to every single image and both learnable in all spatial dimensions, the restriction on a limited number of basis (we use $N=6$ in all of our experiments) can be alleviated.
A key insight of our work is the use of sawtooth coordinate transformation $\gamma(x)=x\bmod{k},k\in\mathbb{R}$, which performs best among tested transformations in image regression tasks. The sawtooth transformation maps input coordinates to a periodic space, implicitly capturing patch-like frequency information as shown in Fig. 2a. However, traditional downsampling in neural networks fails to preserve this periodic structure. To address this, we introduce sawtooth-aware downsampling as in Fig. 2c with our Basis Swin Transformer. This rearrangement allows the network to learn correlations between the periodic features in Sawtooth sampling.
Moreover, in practice, we sample the basis via bilinear or bicubic sampling due to memory constraints, i.e., the feature size is less than image height and width, and this poses a severe problem: These nearby sampled pixel features are actually linear or based on cubic interpolation with respect to the basis, and such inductive bias hinders the representation of non-linearity in images. To resolve this, we propose a modified version, our Factorized Fields:
$\hat{I}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{i}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (3)
where $\psi\in\{\text{sin},\text{cos}\},\alpha_{j}\in R$. Here, inspired by the Fourier Series, with a scalar $\alpha$ which corresponds to an increase in sampling frequency and a transformation function $\psi$ to add to implicit non-linearity functions and modulation of product, our complete Factorized Fields method stands out by the fact that such formulation significantly enhances the ability to represent complex non-linear structures in images and effectively composes high-frequency components between pixels.
Refer to captionRefer to caption
Figure 4: Comparison of Different Representation This figure illustrates the performance comparison on a single-image regression task using three different representations: vanilla factor fields in Eq. 1, our Factorized Fields (Eq. 3), and Plain Fields, which use simple channel-wise tensor concatenation instead of coefficient-basis multiplication. The left panel shows the MAE on the magnitude of different frequency components obtained from the Fourier Transform, while the right panel shows PSNR along training. Although the basis-coefficient decomposition method in [10] yields improvements, it still struggles with slow convergence and inconsistencies between low and high-frequency components. In contrast, by introducing explicit formulations for $\alpha$ and $\psi$, our method aligns different frequency components effectively and achieves superior convergence speed. This advantage becomes especially crucial when the task extends to generalizable applications, such as super-resolution or image compression.
Our Factorized Fields enhance the original Factor Fields [10] in three key ways: (1) non-linear transformation through $\psi\in\{\text{sin},\text{cos}\}$ enables better modeling of complex image patterns, (2) multi-frequency modulation using $\alpha_{j}$ allows explicit control over frequency components, and (3) sawtooth-aware downsampling preserves periodic structure during feature processing. As shown in Fig. 4, our refined representation contributes to large margin of improvement in convergence and performance, demonstrating the modification is the key to superior information modeling.